what hapens to the perimeter of a rectangle if we multiply both its length and widht by 2? by 3? by 10? by n? I really don't know where to start here.
Answer by stanbon(75887)

what hapens to the perimeter of a rectangle if we multiply both its length and widht by 2? by 3? by 10? by n? I really don't know where to start here.
---
Start with the formula:
P = 2(L+W)
---
P/2 = (L+W)
====
Now double length and width:
Multiply both sides by 2 to get:
P = (2L+2W) ; The perimeter doubled
---
Multiply both sides by 3 to get:
3(P/2) = 3L + 3W : The perimeter is multiplied by 3
etc.
Cheers,
Stan H.
============